# On the Intensity-based Inversion Method for Quantitative Quasi-Static Elastography

### v2

@@ 1 Introduction
Figure 1.1 schematically depicts this general workflow in the case of a quasi-static elastography experiment, where the considered sample is imaged before and after deformation, in this case induced by an applied load. As indicated, there are two main approaches for obtaining material parameter estimates following the imaging of the sample, which may itself involve the solution of an inverse problem. The first, and most common one, is known as the two-step approach to quantitative elastography, and involves the solution of two inverse problems: First, in inverse problem 1 (IP1), the displacement or strain inside the sample/tissue undergoing deformation is estimated from its imaging data. Then, in inverse problem 2 (IP2), the material parameters of the sample/tissue are reconstructed from the obtained displacement or strain estimates. IP1 and IP2 are typically considered separately from each other, partly because they belong to quite distinct classes of inverse problems. While IP1 often takes the form of a motion estimation or optical flow problem [65, 66, 3], the material parameter estimation problem IP2 usually falls within the class of PDE-based coefficient-estimation inverse problems, and is in general much better analyzed. Among the vast literature on identifiability of the material parameters, stability, and different reconstruction methods for IP2, see e.g. [8, 9, 10, 13, 14, 19, 20, 25, 31, 33, 37, 38, 51, 56, 57, 75, 47, 43, 35, 17, 29, 6, 2, 5, 4, 12, 36]. Note that many of these works deal with linear elastic material models and time-dependent sample measurements, often inspired by certain applications, but sometimes also for reasons of (numerical) stability and the beneficial uniqueness properties of the time-dependent case. However, in many applications, no dynamic imaging (and hence displacement data) is available and thus, one has to work within the quasi-static setting.
@@
While the two-step approach has generally enjoyed a lot of success in the past, the separate consideration of IP1 and IP2 also has a number of downsides. For example, standard optical flow approaches for IP1 are often “blind” to physical restrictions in the elastography experiment, leading to unrealistic displacement field estimates. This issue, which was recently addressed in [66], also compounds with the issue of non-uniqueness in the material parameter reconstruction problem IP2. Informally, one can say that both IP1 and IP2 may have a large nullspace, while their combination IP1 + IP2 may not. This then motivates so-called one-step approaches to quantitative elastography: instead of sequentially solving IP1 and IP2, one considers the combined inverse problem 3 (IP3), i.e., the direct reconstruction of the material parameters of the sample/tissue undergoing deformation from its imaging data, without an intermediate calculation of the displacement or strain. In this paper, we particularly focus on the intensity-based inversion method (IIM), an intensity-based approach specifically designed for solving the combined IP3. The IIM was introduced in [45] for the specific application of recovering the Young’s modulus of a set of silicone rubber samples imaged with OCT before and after compression. Combining image registration with a linear elasticity-based deformation model, the IIM was able to accurately reconstruct the Young’s modulus for a wide range of sample configurations [45].
@@ 4.2 Convergence analysis II: full noise case
Let Assumption 3.1, 4.1, and 4.3 hold. Then, for $\mathcal{F}$ and $\mathcal{F}^{\delta}$ defined in (7) and (8), respectively, and with $\delta_{\mathcal{F}}=\delta_{\mathcal{F}}(\delta_{\mathcal{I}_{2}},\delta_{G}):=\delta_{\mathcal{I}_{2}}\left|\Omega\right|+´C_{L}\delta_{G}$ there holds
@@
Due to Proposition 4.2 and 4.4, all assumptions of [55, Theorem 2.1] are satisfied, which yields the assertion. Note that while [55] considers finite-dimensional approximations of Tikhonov functionals, the results and proofs themselves also apply in our infinite dimensional setting with noisy operator. ∎
@@
Due to Proposition 4.2 and 4.4, the assertion follows from [55, Theorem 2.3(a)]. ∎
@@ 5.1 General Material Models and Approximate Minimization
for $\mathbf{p}\in\mathcal{P}$, a set of admissible (and potentially function-valued) parameters. As before, our convergence analysis of the IIM (5) remains unaffected by this generalization, as long as $\mathbf{u}$ is uniquely determined by (22) for any admissible parameters $\mathbf{a}\in\mathcal{M}$ and $\mathbf{p}\in\mathcal{P}$, and satisfies all the required (differentiability) assumptions. Furthermore, note that in the full noise case (10), the noise in the deformation $\mathbf{u}^{\delta}$ may be due to noisy parameters $\mathbf{p}^{\delta}$ (such as measured forces), an imprecise material model $\mathcal{L}^{\delta}$, or both.
@@ 5.2 Piecewise constant parameters and convex regularizers
$\boxed{\,\,\,\min\limits_{a_{k}\in\mathcal{M}\subseteq\mathbb{R}^{n}}\;\mathcal{T}_{\alpha,K}^{\delta}(a_{k})\,,\quad\text{where}\quad\mathcal{T}_{\alpha,K}^{\delta}(a_{k}):=\left\|\mathcal{I}_{2}^{\delta}\circ G^{\delta}\left(\sum_{k=1}^{K}a_{k}I_{D_{k}}\right)-\mathcal{I}_{1}^{\delta}\right\|^{2}_{L_{2}(\Omega)}+\alpha\sum_{k=1}^{M}\left|a_{k}\right|^{2}\,.\,\,}$
@@ 6 Application to Linear Elasticity
where $\vec{\boldsymbol{n}}$ is an outward unit normal vector of $\partial\Omega_{1}=\overline{\Gamma_{D}\cup\Gamma_{T}}$, and the stress tensor $\sigma_{\lambda,\mu}$ is defined by
@@
Figure 6.1: Schematic drawing of the object domain $\Omega_{1}$ with displacement-traction boundaries $\Gamma_{D}$ and $\Gamma_{T}$; here $\mathbf{g}_{D}$, is a fixed applied downward displacement.
The linear elasticity model (25) is commonly used in elastography for uniformly isotropic material samples under the assumption of small deformations (infinitesimal strain). The Lamé parameters $\lambda,\mu$ can be directly related to the Young’s modulus $E$ and the Poison ratio $\nu$, which have physical or diagnostic value, depending on the application [76, 69]. Our motivation for considering this linear elastic model are quasi-static optical coherence elastography (OCE) experiments such as those conducted in [45] and schematically depicted in Figure 6.1. There, a rectangular sample is fixed from below and uniformly compressed from above, while the sides of the sample are left to move freely. This corresponds to uniform displacement boundary conditions on the top ($\mathbf{g}_{D}=\text{const.}$) and bottom ($\mathbf{g}_{D}=0$) of the sample, traction free boundary conditions ($\mathbf{g}_{T}=0$) on the sides, and no body forces ($\mathbf{f}=0$). A set of volumetric OCT scans is acquired both before and after compression, which could already be used as the images $\mathcal{I}_{1}^{\delta},\mathcal{I}_{2}^{\delta}$ for the IIM. However, due to the symmetry of the samples considered in [45], the OCT scans are first reduced to 2D maximum intensity projections, which has some practical benefits.
@@
Hence, we have that $\mathcal{M}_{s}(\underline{\mu})\subset H^{s}(\Omega_{1})^{2}\subset L_{\infty}(\Omega_{1})^{2}$, and thus it follows from [35, Theorem 2.1] that for any $(\lambda,\mu)\in\mathcal{M}_{s}(\underline{\mu})$ there exists a unique solution $\tilde{\mathbf{u}}\in V$ of (28) and a constant $c_{\text{LM}}>0$ satisfying
@@
First, note that due to [35, Equation (3.11)], there exists a constant $C>0$ such that
@@
The continuous Fréchet differentiability of $\tilde{\mathbf{u}}$ and the characterization (34) follows directly from [35, Theorem 3.2]. There, the variational form (34) is expressed in the equivalent form
@@
Furthermore, due to [35, Proposition 3.1], there exists constants $C_{1},C_{2}>0$ such that
@@
This follows directly from [52, Theorem 4.16] observing the regularity of $(\lambda,\mu)\in\mathcal{M}_{s}(\underline{\mu})$. ∎
@@
for some $c_{G}=c_{G}(N)$, which is finite since $(h_{\lambda},h_{\mu})\in\mathcal{M}_{s}(\underline{\mu})\subset W^{1,\infty}(\Omega_{1})$ for $s>N/2+1$. Next, we want to apply [52, Theorem 4.16] to obtain the $H^{2}(\Omega)$ regularity of $\mathbf{z}:=\tilde{\mathbf{u}}^{\prime}(\lambda,\mu)(h_{\lambda},h_{\mu})$. For this, it is formally required that $\mathbf{f}_{1}:=\operatorname{div}\left(\sigma_{h_{\lambda},h_{\mu}}(\tilde{\mathbf{u}}(\lambda,\mu)+\Phi)\right)\in L_{2}(\Omega_{1})$. However, a close inspection of the proof shows that $\mathbf{f}_{1}\in L_{2}(\Omega^{\prime})$ is sufficient (cf. also [53, 28, 24]), and we obtain that $\mathbf{z}|_{\Omega}\in H^{2}(\Omega)^{N}$ with
@@
Furthermore, in certain settings of the BVP (25), it is possible to obtain full $H^{2}(\Omega_{1})$ regularity for $\mathbf{u}(\lambda,\mu)$ and thus also for $\mathbf{u}^{\prime}(\lambda,\mu)$. In this case, the existence of the intermediary domain $\Omega^{\prime}$ may be dropped entirely in the above analysis. One example is the case of pure Dirichlet boundary conditions, i.e., $\Gamma_{D}=\partial\Omega$ and $\Gamma_{T}=\emptyset$, for which $H^{2}(\Omega_{1})$ regularity follows if $g_{D}\in H^{\frac{3}{2}}(\Omega_{1})$ and $\partial\Omega_{1}\in C^{1,1}$; cf. [52, Theorem 4.18]. However, for mixed boundary conditions, full $H^{2}(\Omega_{1})$ regularity can typically not be obtained except in very specific circumstances. Nevertheless, if $\Omega$ “touches” $\partial\Omega_{1}$ only on the Dirichlet boundary $\Gamma_{D}$ sufficiently far away from $\Gamma_{T}$, then due to [52, Theorem 4.18] one still has $H^{2}(\Omega)$ regularity assuming $g_{D}\in H^{\frac{3}{2}}(\Gamma_{D})$ and $\partial\Gamma_{D}\in C^{1,1}$, which is sufficient for establishing (12).
@@ 7 Numerical Experiments
In this section, we consider the numerical application of our IIM to the quasi-static linear elastography problem as analyzed in Section 6 on several examples with experimentally motivated simulated data.
@@ 7.1 Problem setting, discretization, and implementation
For the numerical experiments presented below, we consider a rectangular material sample ($N=2$) which is fixed from below and compressed by a certain distance $c_{D}$ from above, while left free to expand on the sides. This setting is motivated by the quasi-static OCE experiments conducted in [45], which due to the symmetry of the considered silicone samples reduce to just such a 2D elastography setting. Now, assuming a linear elastic material, the BVP (25) for the displacement field $\mathbf{u}$ becomes
@@
Figure 7.2: Synthetic OCT images (including speckle) before (left) and after compression (right).
@@ 7.3 Numerical results: noisy images
For the second series of tests, we consider the case of noisy images $\mathcal{I}_{1}^{\delta}$ and $\mathcal{I}_{2}^{\delta}$ on the example of the synthetic OCT sample with a single inclusion; cf. Figure 7.1 (top). To obtain these noisy images, uniformly distributed random noise with different (relative) noise levels $\delta$ is added to the noise-free images $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$. Then, we use the IIM to reconstruct both Lamé parameters $\lambda,\mu$, where in the IIM functional (46) we choose $\alpha=0.1\ast\delta$ as the regularization parameter. Table 7.3 summarizes the corresponding results, each obtained after $100$ iterations of the Nelder-Mead method. It can be seen that the IIM is able to accurately reconstruct both $\lambda,\mu$, with a relative error below $1\%$ even at a noise level of $6\%$. Furthermore, with the exception of the $7\%$ noise case, the relative errors are monotonically decreasing with decreasing noise level, in alignment with our convergence analysis; cf. Figure 7.3 (left).

### v3

@@ 1 Introduction
Figure 1.1 schematically depicts this general workflow in the case of a quasi-static elastography experiment, where the considered sample is imaged before and after deformation, in this case induced by an applied load. As indicated, there are two main approaches for obtaining material parameter estimates following the imaging of the sample, which may itself involve the solution of an inverse problem. The first, and most common one, is known as the two-step approach to quantitative elastography, and involves the solution of two inverse problems: First, in inverse problem 1 (IP1), the displacement or strain inside the sample/tissue undergoing deformation is estimated from its imaging data. Then, in inverse problem 2 (IP2), the material parameters of the sample/tissue are reconstructed from the obtained displacement or strain estimates. IP1 and IP2 are typically considered separately from each other, partly because they belong to quite distinct classes of inverse problems. While IP1 often takes the form of a motion estimation or optical flow problem [77, 78, 3], the material parameter estimation problem IP2 usually falls within the class of PDE-based coefficient-estimation inverse problems, and is in general much better analyzed. Among the vast literature on identifiability of the material parameters, stability, and different reconstruction methods for IP2, see e.g. [8, 9, 10, 13, 14, 19, 20, 25, 31, 36, 40, 41, 61, 66, 67, 89, 53, 49, 38, 17, 29, 6, 2, 5, 4, 12, 39, 55, 56, 57]. Note that many of these works deal with linear elastic material models and time-dependent sample measurements, often inspired by certain applications, but sometimes also for reasons of (numerical) stability and the beneficial uniqueness properties of the time-dependent case. However, in many applications, no dynamic imaging (and hence displacement data) is available and thus, one has to work within the quasi-static setting.
@@
While the two-step approach has generally enjoyed a lot of success in the past, the separate consideration of IP1 and IP2 also has a number of downsides. For example, standard optical flow approaches for IP1 are often “blind” to physical restrictions in the elastography experiment, leading to unrealistic displacement field estimates. This issue, which was recently addressed in [78], also compounds with the issue of non-uniqueness in the material parameter reconstruction problem IP2. Informally, one can say that both IP1 and IP2 may have a large nullspace, while their combination IP1 + IP2 may not. Additionally, since IP1 and IP2 are solved sequentially, data noise and numerical inaccuracies amplify through the decoupled treatment of the corresponding inverse problems. This then motivates so-called one-step approaches to quantitative elastography: instead of sequentially solving IP1 and IP2, one considers the combined inverse problem 3 (IP3), i.e., the direct reconstruction of the material parameters of the sample/tissue undergoing deformation from its imaging data, without an intermediate calculation of the displacement or strain. In this paper, we particularly focus on the intensity-based inversion method (IIM), an intensity-based approach specifically designed for solving the combined IP3. The IIM was introduced in [51] for the specific application of recovering the Young’s modulus of a set of silicone rubber samples imaged with OCT before and after compression. Combining image registration with a linear elasticity-based deformation model, the IIM was able to accurately reconstruct the Young’s modulus for a wide range of sample configurations [51].
@@ 4.2 Convergence analysis II: full noise case
Let Assumption 3.1, 4.1, and 4.3 hold. Then, for $\mathcal{F}$ and $\mathcal{F}^{\delta}$ defined in (7) and (8), respectively, and with $\delta_{\mathcal{F}}=\delta_{\mathcal{F}}(\delta_{\mathcal{I}_{2}},\delta_{G}):=\delta_{\mathcal{I}_{2}}\left|\Omega\right|+C_{L}\delta_{G}$ there holds
@@
Due to Proposition 4.2 and 4.4, all assumptions of [65, Theorem 2.1] are satisfied, which yields the assertion. Note that while [65] considers finite-dimensional approximations of Tikhonov functionals, the results and proofs themselves also apply in our infinite dimensional setting with noisy operator. ∎
@@
Due to Proposition 4.2 and 4.4, the assertion follows from [65, Theorem 2.3(a)]. ∎
@@ 5.1 General Material Models and Approximate Minimization
for $\mathbf{p}\in\mathcal{P}$, a set of admissible (and potentially function-valued) parameters. As before, it can be easily seen from the proofs of the above results that our convergence analysis of the IIM (5) remains unaffected by this generalization, as long as $\mathbf{u}$ is uniquely determined by (22) for any admissible parameters $\mathbf{a}\in\mathcal{M}$ and $\mathbf{p}\in\mathcal{P}$, and satisfies all the required (differentiability) assumptions. Furthermore, note that in the full noise case (10), the noise in the deformation $\mathbf{u}^{\delta}$ may be due to noisy parameters $\mathbf{p}^{\delta}$ (such as measured forces), an imprecise material model $\mathcal{L}^{\delta}$, or both.
@@ 5.2 Piecewise constant parameters and convex regularizers
$\boxed{\,\,\,\min\limits_{a_{k}\in\mathcal{M}\subseteq\mathbb{R}^{n}}\;\mathcal{T}_{\alpha,K}^{\delta}(a_{k})\,,\quad\text{where}\quad\mathcal{T}_{\alpha,K}^{\delta}(a_{k}):=\left\|\mathcal{I}_{2}^{\delta}\circ G^{\delta}\left(\sum_{k=1}^{K}a_{k}\chi_{D_{k}}\right)-\mathcal{I}_{1}^{\delta}\right\|^{2}_{L_{2}(\Omega)}+\alpha\sum_{k=1}^{M}\left|a_{k}\right|^{2}\,.\,\,}$
@@ 6 Application to Linear Elasticity
where $\vec{\boldsymbol{n}}$ is an outward unit normal vector of $\partial\Omega_{1}=\overline{\Gamma_{D}\cup\Gamma_{T}}$, $\mathbf{g}_{D}$ and $\mathbf{g}_{T}$ are (potentially non-constant) boundary functions defined on $\Gamma_{D}$ and $\Gamma_{T}$, respectively, and the stress tensor $\sigma_{\lambda,\mu}$ is defined by
@@
Figure 6.1: Schematic drawing of the object domain $\Omega_{1}$ with displacement-traction boundaries $\Gamma_{D}$ and $\Gamma_{T}$. Here, $\mathbf{g}_{D}$, is a fixed applied downward displacement, which does not have to be constant.
The linear elasticity model (25) is commonly used in elastography for uniformly isotropic material samples under the assumption of small deformations (infinitesimal strain). The Lamé parameters $\lambda,\mu$ can be directly related to the Young’s modulus $E$ and the Poison ratio $\nu$, which have physical or diagnostic value, depending on the application [90, 81]. Our motivation for considering this linear elastic model are quasi-static optical coherence elastography (OCE) experiments such as those conducted in [51] and schematically depicted in Figure 6.1. There, a rectangular sample is fixed from below and uniformly compressed from above, while the sides of the sample are left to move freely. This corresponds to uniform displacement boundary conditions on the top ($\mathbf{g}_{D}=\text{const.}$) and bottom ($\mathbf{g}_{D}=0$) of the sample, traction free boundary conditions ($\mathbf{g}_{T}=0$) on the sides, and no body forces ($\mathbf{f}=0$). (Note that in the subsequent analysis, $\mathbf{g}_{D}$, $\mathbf{g}_{T}$, and $\mathbf{f}$ do not have to be constant but can be functions.) A set of volumetric OCT scans is acquired both before and after compression, which could already be used as the images $\mathcal{I}_{1}^{\delta},\mathcal{I}_{2}^{\delta}$ for the IIM. However, due to the symmetry of the samples considered in [51], the OCT scans are first reduced to 2D maximum intensity projections, which has some practical benefits.
@@
Hence, we have that $\mathcal{M}_{s}(\underline{\mu})\subset H^{s}(\Omega_{1})^{2}\subset L_{\infty}(\Omega_{1})^{2}$, and thus it follows from [38, Theorem 2.1] that for any $(\lambda,\mu)\in\mathcal{M}_{s}(\underline{\mu})$ there exists a unique solution $\tilde{\mathbf{u}}\in V$ of (28) and a constant $c_{\text{LM}}>0$ satisfying
@@
First, note that due to [38, Equation (3.11)], there exists a constant $C>0$ such that
@@
The continuous Fréchet differentiability of $\tilde{\mathbf{u}}$ and the characterization (34) follows directly from [38, Theorem 3.2]. There, the variational form (34) is expressed in the equivalent form
@@
Furthermore, due to [38, Proposition 3.1], there exists constants $C_{1},C_{2}>0$ such that
@@
This follows directly from [62, Theorem 4.16] observing the regularity of $(\lambda,\mu)\in\mathcal{M}_{s}(\underline{\mu})$. ∎
@@
for some $c_{G}=c_{G}(N)$, which is finite since $(h_{\lambda},h_{\mu})\in\mathcal{M}_{s}(\underline{\mu})\subset W^{1,\infty}(\Omega_{1})$ for $s>N/2+1$. Next, we want to apply [62, Theorem 4.16] to obtain the $H^{2}(\Omega)$ regularity of $\mathbf{z}:=\tilde{\mathbf{u}}^{\prime}(\lambda,\mu)(h_{\lambda},h_{\mu})$. For this, it is formally required that $\mathbf{f}_{1}:=\operatorname{div}\left(\sigma_{h_{\lambda},h_{\mu}}(\tilde{\mathbf{u}}(\lambda,\mu)+\Phi)\right)\in L_{2}(\Omega_{1})$. However, a close inspection of the proof shows that $\mathbf{f}_{1}\in L_{2}(\Omega^{\prime})$ is sufficient (cf. also [63, 28, 24]), and we obtain that $\mathbf{z}|_{\Omega}\in H^{2}(\Omega)^{N}$ with
@@
Furthermore, in certain settings of the BVP (25), it is possible to obtain full $H^{2}(\Omega_{1})$ regularity for $\mathbf{u}(\lambda,\mu)$ and thus also for $\mathbf{u}^{\prime}(\lambda,\mu)$. In this case, the existence of the intermediary domain $\Omega^{\prime}$ may be dropped entirely in the above analysis. One example is the case of pure Dirichlet boundary conditions, i.e., $\Gamma_{D}=\partial\Omega$ and $\Gamma_{T}=\emptyset$, for which $H^{2}(\Omega_{1})$ regularity follows if $g_{D}\in H^{\frac{3}{2}}(\Omega_{1})$ and $\partial\Omega_{1}\in C^{1,1}$; cf. [62, Theorem 4.18]. However, for mixed boundary conditions, full $H^{2}(\Omega_{1})$ regularity can typically not be obtained except in very specific circumstances. Nevertheless, if $\Omega$ “touches” $\partial\Omega_{1}$ only on the Dirichlet boundary $\Gamma_{D}$ sufficiently far away from $\Gamma_{T}$, then due to [62, Theorem 4.18] one still has $H^{2}(\Omega)$ regularity assuming $g_{D}\in H^{\frac{3}{2}}(\Gamma_{D})$ and $\partial\Gamma_{D}\in C^{1,1}$, which is sufficient for establishing (12).
@@ 7 Numerical Experiments
In this section, we consider the numerical application of our IIM to the quasi-static linear elastography problem as analyzed in Section 6 on several examples with experimentally motivated simulated data. The considered setting is only one of many possible instances of elastography problems to which our general IIM approach can be applied. As noted above, our particular choice was motivated by the OCE experiments conducted in [51], in which the considered sample has several inclusions whose shape and location can be identified in advance, and where the Lamé parameters are piecewise constant. While we have shown above that the IIM can also be applied to much more general settings (non-constant parameters, unknown inclusion locations, general material models), the numerical setting considered below should be sufficient to numerically demonstrate the theoretical findings of this paper.
@@ 7.1 Problem setting, discretization, and implementation
For the numerical experiments presented below, we consider a rectangular material sample ($N=2$) which is fixed from below and compressed by a certain distance $c_{D}$ from above, while left free to expand on the sides. This setting is motivated by the quasi-static OCE experiments conducted in [51], which due to the symmetry and physical properties of the considered silicone samples reduce to just such a 2D elastography setting. Now, assuming a linear elastic material, the BVP (25) for the displacement field $\mathbf{u}$ becomes
@@
Figure 7.2: Synthetic OCT images (including speckle) before (left) and after compression (right), corresponding to $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$.
@@
Note that real OCT scans do not (necessarily) have a consistent physical unit, but are mainly qualitative intensity images in an arbitrary value range, which are typically rescaled for further processing. This is reflected by our rescaling of the synthetic OCT scans, which are thus unitless intensity images.
@@
Concerning computational costs, note that since in the setting of the considered numerical examples our proposed IIM approach estimates the material parameters via solving the minimization problem (46), the optimization algorithm chosen for this task directly determines the overall computational cost. Similarly, due to the wide array of different two-step approaches, and depending on the methods chosen to implement their two steps, it is difficult to provide a meaningful and comprehensive comparison of their overall computational costs in relation to the IIM. However, note that in most cases, the repeated computation of the displacement field $\mathbf{u}(\mathbf{a})$ for different material parameters $\mathbf{a}$ required in these approaches is typically the numerically most expensive part. For example, the idealized two-step NLI approach introduced in Section 7.6 requires three such function evaluation per iteration, while the Nelder-Mead algorithm used in our implementation of the IIM typically requires only two evaluations per iteration [16]. However, different implementations may require different numbers of function evaluations, and the chosen methods may also require different numbers of iterations. Finally, while the first step in two-step methods does not require any additional evaluations of $\mathbf{u}(\mathbf{a})$, it may still involve a number of computationally expensive calculations depending on the chosen approach.
@@ 7.3 Numerical results: noisy images
For our second series of tests, we consider the case of noisy images $\mathcal{I}_{1}^{\delta}$ and $\mathcal{I}_{2}^{\delta}$ on the example of the synthetic OCT sample with a single inclusion; cf. Figure 7.1 (top). To obtain these noisy images, uniformly distributed random noise with different (relative) noise levels $\delta$ is added to the noise-free images $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$. Then, we use the IIM to reconstruct both Lamé parameters $\lambda,\mu$, where in the IIM functional (46) we choose $\alpha=0.1\ast\delta$ as the regularization parameter, which is consistent with our developed theory. Note that the value $0.1$ was selected by considering the noise level $\delta=1\%$ and testing different choices $\alpha=C\ast\delta$ for a wide range of possible constants $C$, with $C=0.1$ found to be optimal. In practice, different selection criteria such as heuristic parameter choice rules have to be used instead [23, 54, 86, 32, 72, 33, 87, 46, 47, 48, 71, 34, 37].
Table 7.3 summarizes the corresponding results, each obtained after $100$ iterations of the Nelder-Mead method. It can be seen that the IIM is able to accurately reconstruct both $\lambda,\mu$, with a relative error below $1\%$ even at a noise level of $6\%$. Furthermore, with the exception of the $7\%$ noise case, the relative errors are monotonically decreasing with decreasing noise level, in alignment with our convergence analysis; cf. Figure 7.3 (left).